OpenStudy (makayla):

42 is subtracted from 41% of a number. The result is 42. What is the number?

OpenStudy (amistre64):

n(.41) - 42 = 42 n(.41) = 0 n = ??

OpenStudy (amistre64):

ack!! try that again but with brains this time

OpenStudy (amistre64):

n(.41) - 42 = 42 n(.41) = 84 n = 84/.41 = 8400/41

OpenStudy (amistre64):

204.8780487804878048780487804878

OpenStudy (anonymous):

amistre, I can't believe I am not a fan of you yet!! you got it man! :)

OpenStudy (amistre64):

lol ..... thanks :)

OpenStudy (anonymous):

man I supposed right? :P

OpenStudy (makayla):

Amistre64. Thanks for helping me but I made a mistake when posting it should be 42%